OpenStudy (18jonea):

Randex Manufacturing estimates that its weekly profit, P, in hundreds of dollars, can be approximated by the formula P = –2x2 + 4x + 2 where x is the number of units produced per week, in thousands. a. How many units should the company produce per week to earn the maximum profit? b. Find the maximum weekly profit. a. 10 units; b. $1000 a. 100 units; b. $800 a. 1000 units; b. $400 a. 1 unit; b. $500

OpenStudy (18jonea):

@freckles

OpenStudy (freckles):

to earn max profit think about the max of the graph where is the max point located on a parabola? what is that famous v word?

OpenStudy (18jonea):

vertex

OpenStudy (freckles):

yep the parabola is open down because of the negative coefficent on the x^2 part so we definitely have a max and we know max/min occur at the vertex for a parabola

OpenStudy (freckles):

do you know how to find the vertex

OpenStudy (18jonea):

no

OpenStudy (freckles):

do you know how to complete the square?

OpenStudy (freckles):

\[P=-2x^2+4x+2 \\ P=-2(x^2-2x)+2 \\ P=-2(x^2-2x+?-?)+2 \\ P=-2(x^2-2x+?)+(-2)(-?)+2 \\ P=-2(x^2-2x+?)+2?+2\] so x^2-2x+what number will give me a complete square

OpenStudy (18jonea):

100

OpenStudy (18jonea):

@freckles

OpenStudy (18jonea):

is the answer b

OpenStudy (freckles):

how did you get a 100?

OpenStudy (freckles):

actually I guess instead of the choices being labeled a,a,a,a I should imagine it as a,b,c,d :p

OpenStudy (18jonea):

yeah

OpenStudy (18jonea):

and i am just guessing

OpenStudy (freckles):

so you aren't able to answer my one question x^2-2x+what number is going to give me a complete square?

OpenStudy (18jonea):

no how would i do that

OpenStudy (freckles):

\[x^2+kx+(\frac{k}{2})^2=(x+\frac{k}{2})^2\]

OpenStudy (18jonea):

what is k

OpenStudy (18jonea):

@freckles

OpenStudy (freckles):

you compare it to what I gave you to determine k so you can figure out what number to add in to complete the square

OpenStudy (freckles):

\[x^2-2x \text{ is the same as } x^2+kx \text{ when } k=?\]

OpenStudy (18jonea):

2

OpenStudy (freckles):

\[x^2+kx+(\frac{k}{2})^2=(x+\frac{k}{2})^2 \\ \text{ well } k=-2 \text{ not} 2 \\ x^2+kx \text{ is the same as } x^2-2x \text{ when } k=-2 \\ \text{ not just use the formula to complete the \square } \\ x^2+-2x+(\frac{-2}{2})^2=(x+\frac{-2}{2})^2\] you can simplify a bit

OpenStudy (freckles):

\[P=-2x^2+4x+2 \\ P=-2(x^2-2x)+2 \\ P=-2(x^2-2x+?-?)+2 \\ P=-2(x^2-2x+?)+(-2)(-?)+2 \\ P=-2(x^2-2x+?)+2?+2 \\ P=-2(x^2-2x+1)+2(1)+2 \\ P=-2(x-1)^2+2(1)+2\]

OpenStudy (freckles):

this is vertex form \[P=a(x-h)^2+k \text{ where } (h,k) \text{ is vertex }\]

OpenStudy (18jonea):

so its 1

OpenStudy (freckles):

(x,P) where x represents number of units produced each week in thousands and P represents profit in hundreds so when you say 1 what do you mean by 1?

OpenStudy (freckles):

yes the x-coordinate of the vertex is 1 what does that mean though in terms of our problem

OpenStudy (18jonea):

1 unit for 500 ?

OpenStudy (freckles):

are you understanding the x-coordinates are in thousands in the P-coordinate also known as the y-coordinate is in hundreds also how did you get 500?

OpenStudy (freckles):

so the x-coordinate being 1 means you have how many units?

OpenStudy (freckles):

2(1)+2 is 2+2 which is 4 maybe you did 2+1+2 or something not sure.... but the x-coordinate being 1 means? and the y-coordinate being 4 means?

OpenStudy (freckles):

pretend I got the x-coordinate was 2 then that would mean for this problem I have 2000 units is when I get the max profit where the corresponding y value would be the max in hundreds

OpenStudy (freckles):

that is just an example

OpenStudy (freckles):

I'm just asking you to interpret our answer in terms of your problem

OpenStudy (18jonea):

so 1 means 1000

OpenStudy (freckles):

the x-coordinate of the vertex being 1 means we need to produce 1000 units to achieve our max profit

OpenStudy (freckles):

and the y or P-coordinate which is 2(1)+2=4 is in hundreds (dollars)

OpenStudy (freckles):

so what is our max profit ?

OpenStudy (18jonea):

so 1000 units 400$

OpenStudy (freckles):

yes producing 1000 units will achieve our max profit of 400 dollars

OpenStudy (18jonea):

@freckles

OpenStudy (freckles):

yes?

OpenStudy (18jonea):

?

OpenStudy (freckles):

do you have another question?

OpenStudy (freckles):

if you posted another question I really do not see it...

OpenStudy (18jonea):

no i'm all set thank you so much

OpenStudy (freckles):

ok I was confused